OpenStudy (anonymous):

A does a work in 36 days,B does it in 24 days and C does it in 96 days.They started doing work together,but A left the work after 9 days ,B left it before 9 days of its compeletion,how long A,B and C work together to complete the work.

OpenStudy (anonymous):

work by A in one day = 1/36 work by B in one day = 1/24 work by C in one day = 1/96

ganeshie8 (ganeshie8):

9 days

OpenStudy (anonymous):

in first 9 days A,B,C work together total work in 9 days = 9\16

OpenStudy (anonymous):

as on last nine days C has to do the work alone therefore work on last 9 days = 9/96

ganeshie8 (ganeshie8):

\(\frac{9}{16} + (\frac{1}{24} + \frac{1}{96})d + \frac{9}{96} = 1\)

OpenStudy (anonymous):

work remaining is 1 - 9/16 - 9/96 now this work can be done only B and C

ganeshie8 (ganeshie8):

total days it took for them to finish the work = 9 + d + 9